Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 3.3 pb-1.

Monday December 1
The day shift began with the Tevatron (TeV) recovering from a quench, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 10:33 AM, Pbar experts began a reverse proton study period.

A Linac RF station (KRF2) started sparking around 1:30 PM. A Linac technician worked on the station for an hour, but reported slow progress. The station came back online at 3:30 PM.

A Pbar sextupole (SEX12) power supply tripped off at 2:41 PM and caused the loss of some antiprotons. An expert investigated.

Operations put the TeV through a dry squeeze at 3:32 PM.

MiniBooNE experts accessed their tunnel.

TeV experts had been conducting studies all day. Unfortunately, the TeV quenched at 4:36 PM during a study.

Pbar began stacking at 5 PM.

At 5:39 PM, two Booster RF stations (BRF3, 4, 5 & 6) tripped off due to a vacuum burst. They reset okay.

At 11:19 PM, the TeV suffered a sector A1 quench during an abort.

Tuesday December 2
The midnight shift began with the TeV recovering from a quench, with Pbar stacking, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 2:40 AM.

At 5:12 AM, the TeV quenched at sector F3.

At 6 AM, operators reported having problems with a Linac RF station (LRF1).

Operators prepared D-zero (D0) for an access at 7:30 AM. The experimenters were out of access by 9 AM.

Operations put the TeV through a dry squeeze at 9:31 AM.

Pbar stopped stacking at 9:47 AM to conduct studies.

Operations loaded the TeV with a 12x0 study-store at 11 AM.

A Booster expert began aperture scans at 1:12 PM. This study shut off beam to MiniBooNE.

At 4:30 PM, Pbar studies ended and stacking resumed.

Booster studies came to a point where Operations could also send beam to MiniBooNE at 4:34 PM.

Wednesday December 3
The midnight shift began with many studies occurring, with Pbar stacking, and with MiniBooNE taking beam.

At 12:21 AM, the TeV suffered a sector B1 quench.

Around 5 AM, TeV experts began a quench-at-abort study.

At 7:48 AM, Booster achieved its first 5E12 pulse of beam since before the shutdown.

Operations put the TeV through a dry squeeze at 7:52 AM.

Today's Plans
The TeV experts will continue their quench-at-abort studies. The Recycler will take shots of antiprotons from Pbar sometime during the day.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/3/2003   email Fermilab